OpenStudy (anonymous):

if a temperature is 140 degrees Fahrenheit then what would it be in celcius

OpenStudy (wolf1728):

Celsius = (F-32) x 5/9 Celsius = (140-32) x 5/9 Celsius = 108 x 5/9 = 60° Celsius Formula is here: http://www.1728.org/convtmp2.htm